OpenStudy (owlsarecool):

Help plz i was sick on the day we learned about it :( Solve each system of equations. 2x+y=11 y=x-1

OpenStudy (caozeyuan):

ok, so 2x+y=11 and y-x=-1

OpenStudy (steve816):

I would say use substitution. Plug in x-1 to y in the first equation.

OpenStudy (caozeyuan):

now 3x=11-(-1)=12

OpenStudy (caozeyuan):

and x=4

OpenStudy (caozeyuan):

so y=3

OpenStudy (owlsarecool):

I am a little confused can you you that together for me? i can solve but i need to know what i am solving and i am over all confused. sorry i am trying :(

OpenStudy (caozeyuan):

basically if trying to cancel the y in the two equations

OpenStudy (caozeyuan):

so that the result is a single equation with only x in it

OpenStudy (owlsarecool):

ok...

OpenStudy (caozeyuan):

to do that, I subtract the left side of second euqatoon from the left side of the first

OpenStudy (caozeyuan):

it has to equal to the result of subtracting the right side of the second from right side of the first

OpenStudy (caozeyuan):

now the left of new euqation is 2x, agree?

OpenStudy (caozeyuan):

right of new queation is 11-x+1

OpenStudy (caozeyuan):

rearaange gives you 3x=12

zepdrix (zepdrix):

haha, owls are cool c: especially blue owls!

OpenStudy (steve816):

@steve816 is the coolest fella I've ever met!

OpenStudy (owlsarecool):

yes but i still do not see exactly how that is shown.. like 2(x+y)=11?

OpenStudy (owlsarecool):

thats not right is it :( i am so confused im sorry i am trying but i still dont get it

OpenStudy (caozeyuan):

what is 2x+y-y?

OpenStudy (owlsarecool):

2x? Sorrry if i take forever to respond its my computer is not working right

OpenStudy (caozeyuan):

good, now how about 11-(x-1)?

OpenStudy (owlsarecool):

12x? i do not think i did that right...

OpenStudy (caozeyuan):

ok, so you should try again

OpenStudy (caozeyuan):

where exactly did you strcuk?

OpenStudy (owlsarecool):

um never mind i figured it out is the answer (4,3)

OpenStudy (caozeyuan):

great

OpenStudy (owlsarecool):

Yay

OpenStudy (owlsarecool):

thanks